OpenStudy (abbycross167):

Can someone please help me on this math question? please? I'll fan and give a medal!!

OpenStudy (abbycross167):

For questions 17 through 20 use f(x) = x2 – x + 2 17. find f(2)

OpenStudy (abbycross167):

@Abbles

OpenStudy (abbycross167):

would it be f(x) = x2 – x + 2 --- f(2) = (2)2 – (2) + 2 or -- f(2) = x2 – (2) + 2 ?

OpenStudy (abbles):

Assuming x2 means x squared, it would look like: f(2) = 2^2 - 2 + 2 So you're first one looks about right :)

OpenStudy (abbycross167):

okay I'll do that, but the question didn't have x^2 just x2

OpenStudy (abbycross167):

My teacher might have forgotten to square it... I guess

OpenStudy (sshayer):

just plug x=2 and find the solution.

OpenStudy (sshayer):

can you show me the solution?

OpenStudy (sshayer):

\[f(x)=x^2-x+2\] \[f(2)=2^2-2+2=?\]

OpenStudy (abbles):

If it is indeed squared, the solution would be: f(2) = 2^2 - 2 + 2 f(2) = 4 - 2 + 2 f(2) = 4

OpenStudy (abbles):

Unless the 2 is actually before the x on the first term? In which case it would be: f(2) = 2(2) - 2 + 2 f(2) = 4 - 2 + 2 f(2) = 4 Which actually turns out to be the same thing!

OpenStudy (sshayer):

correct.

OpenStudy (abbycross167):

@Abbles Thank you!!!! You don't know how much I appreciate your help!

OpenStudy (sshayer):

no, when written x2-x+2 it is \[x^2-x+2\] when written 2x-x+2 it is 2*x-x+2

OpenStudy (abbles):

Your welcome :) ask away if you have any more Qs, I'm happy to help.

OpenStudy (abbycross167):

Could you check this answer? please? 18. Find f(0) f(0)= 0^2 – 0 + 2 f(0)= 0 – 0 +2 f(0)=2

OpenStudy (sshayer):

correct.

OpenStudy (abbycross167):

f(x) = x^2 – x + 2 19. Find f(-4) f(-4)= -4^2 – (-4) + 2 f(-4)= -16-(-4)+2 f(-4)= -12+2 f(-4)= -10

OpenStudy (abbles):

Your f(0) is correct but not the f(4).

OpenStudy (abbles):

What is -4 times -4?

OpenStudy (sshayer):

correction \[f(-4)=(-4)^2-(-4)+2=?\]

OpenStudy (abbycross167):

would it be 16?

OpenStudy (abbles):

Yes. f(x) = x^2 – x + 2 f(-4)=( -4)^2 – (-4) + 2 f(-4)= 16 + 4 +2 f(-4)= 22

OpenStudy (abbycross167):

Thank you!! Could you check this last problem? Please? f(x) = x^2 – x + 2 20. Find f(1) f(1)= 1^2 – 1 + 2 f(1)=1-1+2 f(1)=2

OpenStudy (abbles):

Sure thing. Looks all good :)

OpenStudy (abbycross167):

Great! thank you for everything!! You're awesome!! :)